OpenStudy (anonymous):

solve by using elimination 3x+7y=1 2x+9y=5

OpenStudy (anonymous):

think it is (-2,1)

OpenStudy (anonymous):

Elimination works by adding both equations together and eliminating one variable in the process. So first you need for one of the variables to have the same coefficient on both equations, except with different signs. Step 1: Multiply the first equation times 2 and the second one times 3 6x+14y=2 6x+27y=15 Step 2: Subtract the first equation from the second 13y=13 Step 3: Solve for the remaining variable y=1 Step 4: Substitute the new value for that variable into either equation and solve for the other variable. 3x+7y=1 3x+7*(1)+1 6x+7=1 6x=1-7 6x=-6 x=-1